OpenStudy (anonymous):

More integration by substitution...problem below:

OpenStudy (anonymous):

\[\int\limits (x^2 +1)^3 dx\]I tried u=x^2+1...no dice.

OpenStudy (anonymous):

Is there an efficient way to go about this or do I actually have to multiply it out first?

zepdrix (zepdrix):

No you certainly don't have to multiply it out :) That is definitely one option though, and it wouldn't be too difficult. Alternatively you can do a trig substitution. :D Might actually be more work though on this particular one ^^ Depends what you wanna try c:

OpenStudy (anonymous):

Ah, thought it might be trig...just watched videos on that today...

zepdrix (zepdrix):

For trig subs you really want to have these 2 identities in the front of your mind, so you can easily recognize them when things like this show up :D \[\sin^2x+\cos^2x=1\]\[\tan^2x+1=\sec^2x\] Think you can get through it ok? :D

OpenStudy (anonymous):

Ah, yes, thank you @zepdrix. I can see where some of the gaps in my self-instruction are. Looks like I'm going to need to drill trig identities.

OpenStudy (anonymous):

So it looks like I'd do this: \[u=\sec^2x\] Well, du is going to be duh for me for the moment. :) Thanks, though. I know what I need to study now.

zepdrix (zepdrix):

Ah :)

zepdrix (zepdrix):

 :O